OpenStudy (anonymous):

Can anyone help me prove (7csc(X)-7)/cot(x)=(7cot(x))/(csc(x)+1)

OpenStudy (michele_laino):

I can rewrite the left side like below: \[\Large \frac{{7\csc x - 7}}{{\cot x}} = \frac{{\frac{7}{{\sin x}} - 7}}{{\frac{{\cos x}}{{\sin x}}}}\]

OpenStudy (michele_laino):

is it ok?

OpenStudy (anonymous):

Yes

OpenStudy (michele_laino):

now, we have: \[\Large \frac{7}{{\sin x}} - 7 = \frac{{7 - 7\sin x}}{{\sin x}}\]

OpenStudy (anonymous):

Now I'm lost. Where did the cos(x) go?

OpenStudy (michele_laino):

that is the numerator only

OpenStudy (anonymous):

Okay that makes more sense

OpenStudy (michele_laino):

so, we can write: \[\Large \begin{gathered} \frac{{\frac{7}{{\sin x}} - 7}}{{\frac{{\cos x}}{{\sin x}}}} = \left( {\frac{7}{{\sin x}} - 7} \right) \times \frac{{\sin x}}{{\cos x}} = \hfill \\ \hfill \\ = \frac{{7 - 7\sin x}}{{\sin x}} \times \frac{{\sin x}}{{\cos x}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

\[\frac{ 7-7sinx }{ cosx } ?\]

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

and we can rewrite it as below: \[\Large \frac{{7\left( {1 - \sin x} \right)}}{{\sin x}}\]

OpenStudy (michele_laino):

is it ok?

OpenStudy (anonymous):

Yes

OpenStudy (michele_laino):

oops..the right expression is: \[\frac{{7\left( {1 - \sin x} \right)}}{{\cos x}}\]

OpenStudy (michele_laino):

Now, I multiply numerator and denominator by cosx, so I get: \[\Large \frac{{7\left( {1 - \sin x} \right)}}{{\cos x}} \times \frac{{\cos x}}{{\cos x}} = \frac{{7\left( {1 - \sin x} \right)\cos x}}{{{{\left( {\cos x} \right)}^2}}}\]

OpenStudy (anonymous):

Then we can use the Pyth. ID right?

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

and we get this: \[\Large \begin{gathered} \frac{{7\left( {1 - \sin x} \right)\cos x}}{{{{\left( {\cos x} \right)}^2}}} = \frac{{7\left( {1 - \sin x} \right)\cos x}}{{1 - {{\left( {\sin x} \right)}^2}}} = \hfill \\ \hfill \\ = \frac{{7\left( {1 - \sin x} \right)\cos x}}{{\left( {1 - \sin x} \right)\left( {1 + \sin x} \right)}} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

Then the 1-sinx cancel out

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

you should get this: \[\Large \frac{{7\left( {1 - \sin x} \right)\cos x}}{{\left( {1 - \sin x} \right)\left( {1 + \sin x} \right)}} = \frac{{7\cos x}}{{1 + \sin x}}\]

OpenStudy (anonymous):

Yes

OpenStudy (michele_laino):

ok! Now I divide both numerator and denominator by sinx, and I get: \[\Large \frac{{7\cos x}}{{1 + \sin x}} = \frac{{7\frac{{\cos x}}{{\sin x}}}}{{\frac{{1 + \sin x}}{{\sin x}}}}\]

OpenStudy (michele_laino):

ok?

OpenStudy (anonymous):

So the numerator would be 7cotx

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

whereas the denominator can be rewritten as below:

OpenStudy (michele_laino):

\[\Large \begin{gathered} \frac{{1 + \sin x}}{{\sin x}} = \frac{1}{{\sin x}} + \frac{{\sin x}}{{\sin x}} = \hfill \\ \hfill \\ = \frac{1}{{\sin x}} + 1 = \csc x + 1 \hfill \\ \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

That's what I got. Thank you so much

OpenStudy (michele_laino):

Thank you!

OpenStudy (anonymous):

by definitions 1/sin x= csc x cos x/sin x = cot x so [(7/sinx) -7]/(cosx/sin x)= (7csc x -7)/cot (x)